What sample size should be obtained to determine a 94% confidence within 4%?




1
Expert's answer
2021-03-08T19:13:17-0500

Sample size formula:

"n = \frac{Z_{α/2}^2 pq}{E^2}"

P unknown:

"n = \frac{Z_{α/2}^2 0.25}{E^2}"

Determine "Z_{\u03b1\/2}" for confidence level of 94% using table:

"Z_{0.03} = 1.89"

"n = \frac{1.89^2 \times 0.25}{0.04^2} = \frac{0.893}{0.0016} = 558"

Answer: 558